OpenStudy (anonymous):

Express answers in simplest exact form. What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot? A =? PI sq. ft.

OpenStudy (anonymous):

area = pi *r^2 r=1foot so area of 1/4 the circle (360 degrees in a circle your asked for 90) answer: pi/4

OpenStudy (anonymous):

heY THanks i juss need the answer im in a huge rush

OpenStudy (anonymous):

i got 12.56

OpenStudy (anonymous):

\[\pi/4\]=0.785 ft.square

OpenStudy (anonymous):

it saYs its rong